OpenStudy (anonymous):

Solve the system of equations. x + 3y = −1 2x + 2y = 6 (−4, 1) (2, −1) (4, −1) (5, −2)

OpenStudy (anonymous):

@Michele_Laino can u help me

OpenStudy (mathstudent55):

Do you have a preferred method, substitution, addition, etc.?

OpenStudy (anonymous):

Use the substitution method and isolate x in the first equation. x + 3y = -1 x + 3y - 3y = -1 - 3y x = -1 - 3y After you've isolated x, plug in what x equals onto the second equation. 2x + 2y = 6 2(-1 - 3y) + 2y = 6 Distribute the expression in parenthesis. Next, combine all like terms and solve for y. If you've received y, just plug it back in the first original equation to receive the value of x. That will be your (x, y) value, or your solution.

OpenStudy (anonymous):

Type both equations in this link. It will show you the steps. http://www.mathpapa.com/calc.html?q=x+y%3D7%2C%20x+2y%3D11

OpenStudy (anonymous):

@Q345